OpenStudy (anonymous):

find the coordinate of the midpoint of the segment S (4,-1) and T(6,0)

OpenStudy (lgbasallote):

\[\huge \text{midpoint} = \left( \frac{x_1 + x_2}{2} , \; \frac{y_1 + y_2}{2}\right)\] does that help?

OpenStudy (lgbasallote):

so in this case.. \[\huge \text{m.p.} = \left(\frac{4 + 6}{2} , \; \frac{-1 + 0}{2}\right)\] got it?

OpenStudy (anonymous):

so \[\ \frac{ 10 }{ 2} \frac{ 0 }{ 2 } \] am i right?

OpenStudy (lgbasallote):

no... \[\left(\frac{10}2 \; , \; -\frac 12 \right)\] -1 + 0 = -1

OpenStudy (anonymous):

so \[\left( \frac{ 9 }{ 2 } \right) ?\]